OpenStudy (anonymous):

Find the 8th term of the arithmetic sequence: 3, 8, 13, 18, ... Find the 11th term of the arithmetic sequence: -3, -5, -7, -9, ...

OpenStudy (tad1):

difference is 5 3 + 7*5 = 38

OpenStudy (tad1):

11th term difference is -2 -3 +10(-2) = -23

OpenStudy (amistre64):

d = -2 an = a1 + (n-1)(d) a(11) = -3 + (10)(-2) = -23; yep

OpenStudy (anonymous):

Find the indicated sum for the arithmetic series: \[\sum_{k-1}^{15} (50-20k)\]

OpenStudy (anonymous):

last one please

OpenStudy (amistre64):

the sum from 1 to 15 eh ....... i can do it the long way but I think the short cut is: (a1 + a15)(15) ------------- 2 ill have to dbl chk meself :)

OpenStudy (anonymous):

the main answer is 2?

OpenStudy (amistre64):

just hold on :)

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

im thinking if I did it right its gonna be -110; but let me check on that first

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

i guess that wasnt the formula I was thinking it was :) When I add all the values up I get -1650

OpenStudy (anonymous):

ok ya thats right thanks :)

OpenStudy (amistre64):

i had the right formula i just forgot to multiply in the 15 :) same answer